The value of a boat is depreciating at a rate of 7% per hour. In 2004, the boat was worth $192,000. Find the value of the boat in 2013.
Answer by Cromlix(4381)

Hi there,
Depreciating at 7%
100 - 7 = 93%
93% = 93/100 = 0.93
0.93 is your factor of depreciation
There are 9 years between 2004 and 2013
So, you set 9 as a power on your factor
0.93^9
Multiply your original value
$192,000 x 0.93^9
= $99918.93
This is the value of the boat in 2013.
Hope this helps :-)